AN ENGINE RUNNING AT 5000rpm produces a power output of 210kw.the mean radius of clutch is 15cm .if the coeficient of friction between surfaces is 0.35 and a clamping force of 20kn,calculate the torque transmitted by the clutch
1
Expert's answer
2020-09-17T12:32:38-0400
"torque = force \times coef \times mean radius"

radius = 15 cm = 0.15m

coef = 0.35

clamping force = 20kn "= 20 \times10 ^{3}"

"torque = 20 \times 10 ^{3} \times 0.35 \times 0.15"

= 1050 NM